OpenStudy (anonymous):

Name of polyhedral Number of faces Number of vertices Number of edges Solid 1 10 x 24 Solid 2 y 7 12 What is the value of x and y? x =16, y = 5 x =14, y = 7 x =7, y = 16 x =16, y =7

OpenStudy (anonymous):

I just want the answers plz I did the rest by my self

OpenStudy (anonymous):

Use Euler's formula.

OpenStudy (anonymous):

can u plz tell me what that is or show me what it is

OpenStudy (anonymous):

V - E + F = 2 V Vertices E Edges F Faces

OpenStudy (anonymous):

ok thanks so much

OpenStudy (anonymous):

wait what do I do with the 2

OpenStudy (anonymous):

It's an equation Find the unknown.

OpenStudy (anonymous):

could u make up a question and then answer it so I can see how to do it

OpenStudy (anonymous):

I'll solve one for you. Solid 1: F = 10. E =24. V = x. x - 24 +10 = 2. x - 14 = 2. x = 16.

OpenStudy (anonymous):

ok I get that but the solid2 is not working out for me I keep getting 17

OpenStudy (anonymous):

-17

OpenStudy (anonymous):

Show your work.

OpenStudy (anonymous):

F=y V=7 E=12 y+7-12=2 y-5=2 y=7

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

thanks so much